The website functions.wolfram.com

Oleg Marichev Wolfran Research Inc.

Currently the website www.functions.wolfram.com includes more than 37000 formulas for about 250 elementary and special functions. The structure and organization of the website will be discussed in detail. An overview is given how to manage, maintain, develop, and update this large mathematical knowledge base and how to deliver various formats types (Mathematica notebooks, html, MathML, and pdf). Future development directions will be outlined. New formulas like new connection between Sin[z] and Sqrt[1-Cos[z]ˆ 2] will be demonstrated.